OpenStudy (anonymous):

What is the approximate distance between Portland, Oregon (45° north, 122° west), and San Francisco, California (37° north, 122° west). Use 3,960 miles for the radius of Earth. 258 miles 553 miles 720 miles 1,346 miles

OpenStudy (anonymous):

@uChezzy

OpenStudy (leonardo0430):

720 miles :)

OpenStudy (anonymous):

pinky promise?

OpenStudy (leonardo0430):

pony promise ;)

OpenStudy (campbell_st):

oops the sector angle is 8 degrees on a great circle so its \[d = \frac{8}{360} \times 2 \times \pi \times r\]

OpenStudy (campbell_st):

just check it with the formula... as 720 is not correct

OpenStudy (anonymous):

what is correct???

OpenStudy (anonymous):

@campbell_st